# Deeply virtual $\phi$-meson production near threshold

### v1

@@ V.2 SoLID
At JLab, the future SoLID program in Hall A provides the luminosity, particle identification, and acceptance needed to perform precise measurements of rare processes. An experiment to study electroproduction of $\phi$ mesons in SoLID could run concurrently with the already approved SoLID $J/\psi$ experiment. The $J/\psi$ experiment plans to take 50 days of physics running with an 11 GeV $e^{-}$ beam at a current of 3 $\mu$A on a 15cm long LH2 target, producing an instantaneous luminosity of approximately 1037 cm-2/s and collecting an integrated luminosity of 43.2 ab-1. To evaluate SoLID’s capabilities for this measurement, we simulated 43.2 ab-1 of $Q^{2}>2$ GeV2 events using lAger at a beam energy of 10.6 GeV. A comprehensive summary of the SoLID detector is provided in Ref. Arrington et al. (2023).
@@
The resolutions and detector acceptances assumed in this study were determined by the SoLID collaboration with a dedicated GEANT4 simulation in the GEMC framework GEMC (started 2007); libsolgem (started 2011); SOLID_GEMC (started 2010); SoLIDTracking (started 2016); Allison et al. (2016). We make use of these results in a fast MC to smear the four-vectors of generated particles in accordance with the expected resolutions. The acceptance of the SoLID $J/\psi$ configuration in momentum and polar angle $\theta$ is shown in Fig. 13. The fast MC queries the acceptance map and randomly discards or accepts particles based on their momentum and polar angle in accordance with the expected acceptance. Compared to the acceptance map determined with the SoLID GEANT4 simulation, for the purposes of making realistic projections we scale the acceptance by a factor of 0.9 to simulate tracking inefficiencies in the high-rate environment at SoLID. We furthermore only consider events in which the proton and kaons enter the FAD, such that they can be efficiently identified by the particle identification detectors. We assume perfect identification of kaons and protons in the FAD, which should be a reasonable approximation since background pions can be rejected by the heavy gas Cherenkov and protons can be identified by the time-of-flight.
@@
Figure 13: The fractional acceptance of SoLID for charged particles in momentum and polar angle. Left: Full acceptance for particles of momenta up to 11 GeV. Right: Zoomed acceptance for particles in the momentum range particularly relevant for reactions at low values of $|t|$.
Figure 14: Momentum (Left) and polar angle (Right) distributions for particles in fully reconstructed events with $41.96<W<$ 4 GeV and $Q^{2}>2$ GeV2. The gap between the forward-angle and large-angle detectors is visible around 16 degrees.
@@
For comparison with the results for the EIC, in the range of $2.82.4<W<$ 2.8 GeV and $Q^{2}>3$ GeV2, we project that SoLID will reconstruct around 270,000 events with all four final-state particles being detected. The SoLID acceptance in the near-threshold region ($W<2.8$ GeV) depends strongly on $Q^{2}$, ranging from almost 10% for events with $87<Q^{2}<$ 8 GeV2 to around 0.2% for $43<Q^{2}<$ 4 GeV2. Similarly to the EIC, the kinematic reach and event statistics can be improved beyond what is presented here if only three of the four final-state particles are required to be detected. Another possible way to improve the statistics is to reconstruct the final state $e^{\prime},p,$ and $K_{s}^{0}$ from the $\phi\rightarrow K_{s}^{0}K^{0}_{L}$ decay and infer the $K^{0}_{L}$ from missing mass. As in Fig. 11, the theory curves are generated at the bin centers in $Q^{2}$ and $W$. Due to the form factor of (62), the highest statistics are available at low values of $|t|$. The region $|t|\lesssim Q^{2}/3$ is observed to be well-populated with statistically precise data points. The acceptance of SoLID increases at higher $Q^{2}$, due to the increased likelihood that the scattered proton and the $\phi$ decay products have enough transverse momentum to enter the FAD. This effect produces the unintuitive improvement in the number of measurable data points and uncertainties at higher $Q^{2}$, in spite of the fact that the cross section falls rapidly as a function of $Q$. As can be seen from Fig. 15 and Fig. 1, $\xi>0.4$ for values of $W\lesssim 2.6$ at $Q^{2}=6$ GeV2. Additionally, the range of $W$ for which $\xi>0.4$ grows at higher $Q^{2}$. Therefore, the region $|t|\lesssim\frac{Q^{2}}{3}$ and $\xi>0.4$ should be experimentally accessible with a reasonable degree of precision at SoLID.
@@ VI Conclusion
In this work, we have presented a new perturbative QCD analysis of near-threshold $\phi$-meson electroproduction in $ep$ scattering. The biggest advantage of being near the threshold is that the skewness variable $\xi$ is order unity. More specifically, we have identified the range $0.60.4<\xi<$ 0.6 as the practical region of interest. The DVMP amplitudes are then dominated by the gravitational form factors Hatta and Strikman (2021); Guo et al. (2021). This can be seen most dramatically in the conformal partial wave expansion where keeping only the $j=1$ moment provides a very good approximation to the entire amplitude Guo et al. (2024). We have demonstrated, for the first time, that this ‘threshold approximation’ works also at next-to-leading order in perturbation theory and even in the quark channel. Using the GK model as an example, we estimate that the truncation error can be as good as 5% (10%) at the amplitude (cross section) level in the kinematics considered in this paper. This is expected to be smaller than other sources of errors such as higher twist corrections. In principle, the approximation should work even better at higher $Q^{2}$, although measurements then become challenging due to the fast falloff of the cross section with increasing $Q^{2}$.

### v3

@@ V.2 SoLID
[…]
@@
The resolutions and detector acceptances assumed in this study were determined by the SoLID collaboration with a dedicated GEANT4 simulation in the GEMC framework Ungaro (started 2007); Xiong (started 2011); Collaboration (started 2010); Xiong (started 2016); Allison et al. (2016). We make use of these results in a fast MC to smear the four-vectors of generated particles in accordance with the expected resolutions. The single-particle acceptance of the SoLID $J/\psi$ configuration in momentum and polar angle $\theta$ is shown in Fig. 13. The distributions in momentum and polar angle for electrons, kaons, and protons in events where all four final-state particles are reconstructed are provided in Fig. 14. The fast MC queries the acceptance map and randomly discards or accepts particles based on their momentum and polar angle in accordance with the expected acceptance. Compared to the acceptance map determined with the SoLID GEANT4 simulation, for the purposes of making realistic projections we scale the acceptance by a factor of 0.9 to simulate tracking inefficiencies in the high-rate environment at SoLID. We furthermore only consider events in which the proton and kaons enter the FAD, such that they can be efficiently identified by the particle identification detectors. We assume perfect identification of kaons and protons in the FAD, which should be a reasonable approximation since background pions can be rejected by the heavy gas Cherenkov and protons can be identified by the time-of-flight.
@@
Figure 13: The fractional acceptance of SoLID for charged particles in momentum and polar angle. Left: Full acceptance for particles of momenta up to $1$ GeV. Right: Zoomed acceptance for particles in the momentum range particularly relevant for reactions at low values of $|t|$.
Figure 14: Momentum (Left) and polar angle (Right) distributions for particles in fully reconstructed events with $1.96<W<4$ GeV and $Q^{2}>2$ GeV2. The gap between the forward-angle and large-angle detectors is visible around 16 degrees.
@@
For comparison with the results for the EIC, in the range of $2.4<W<2.8$ GeV and $Q^{2}>3$ GeV2, we project that SoLID will reconstruct around 270,000 events with all four final-state particles being detected. The SoLID acceptance in the near-threshold region ($W<2.8$ GeV) depends strongly on $Q^{2}$, ranging from almost 10% for events with $7<Q^{2}<8$ GeV2 to around 0.2% for $3<Q^{2}<4$ GeV2. Similarly to the EIC, the kinematic reach and event statistics can be improved beyond what is presented here if only three of the four final-state particles are required to be detected. Another possible way to improve the statistics is to reconstruct the final state $e^{\prime},p,$ and $K_{s}^{0}$ from the $\phi\rightarrow K_{s}^{0}K^{0}_{L}$ decay and infer the $K^{0}_{L}$ from missing mass. As in Fig. 11, the theory curves are generated at the bin centers in $Q^{2}$ and $W$. Due to the form factor of (62), the highest statistics are available at low values of $|t|$. The region $|t|\lesssim Q^{2}/3$ is observed to be well-populated with statistically precise data points. The acceptance of SoLID increases at higher $Q^{2}$, due to the increased likelihood that the scattered proton and the $\phi$ decay products have enough transverse momentum to enter the FAD. This effect produces the unintuitive improvement in the number of measurable data points and uncertainties at higher $Q^{2}$, in spite of the fact that the cross section falls rapidly as a function of $Q$. As can be seen from Fig. 15 and Fig. 1, $\xi>0.4$ for values of $W\lesssim 2.6$ at $Q^{2}=6$ GeV2. Additionally, the range of $W$ for which $\xi>0.4$ grows at higher $Q^{2}$. Therefore, the region $|t|\lesssim\frac{Q^{2}}{3}$ and $\xi>0.4$ should be experimentally accessible with a reasonable degree of precision at SoLID.
@@ VI Conclusion
In this work, we have presented a new perturbative QCD analysis of near-threshold $\phi$-meson electroproduction in $ep$ scattering. The biggest advantage of being near the threshold is that the skewness variable $\xi$ is order unity. More specifically, we have identified the range $0.4<\xi<0.6$ as the practical region of interest. The DVMP amplitudes are then dominated by the gravitational form factors Hatta and Strikman (2021); Guo et al. (2021). This can be seen most dramatically in the conformal partial wave expansion where keeping only the $j=1$ moment provides a very good approximation to the entire amplitude Guo et al. (2024). We have demonstrated, for the first time, that this ‘threshold approximation’ works also at next-to-leading order in perturbation theory and even in the quark channel. Using the GK model as an example, we estimate that the truncation error can be as good as 5% (10%) at the amplitude (cross section) level in the kinematics considered in this paper. This is expected to be smaller than other sources of errors such as higher twist corrections. In principle, the approximation should work even better at higher $Q^{2}$, although measurements then become challenging due to the fast falloff of the cross section with increasing $Q^{2}$.
@@
Note added
On the same day our paper was posted on arXiv, a preprint Guo et al. (2025) was released, where the authors discuss the connection between $J/\psi$ photoproduction near the threshold and the gluon GFFs to next-to-leading order. More recently, the validity of the threshold approximation has been tested also for the pion target Hatta and Schoenleber (2025).